OpenStudy (anonymous):

Mrs. West is 14 years younger than her aunt. If Mrs. West's age in years is as much below 60 as her aunt's age is over 40, how old is each?

OpenStudy (anonymous):

17

OpenStudy (anonymous):

Let's say aunt's age =x, that means Mrs West age =x-14. The same time the age in years for aunt is 40+y (where y is the number of years she is above 40). the age of Mrs West in years is 60-y ( y is the number of years she is below 60, coz both numbers are same) now we get an equation. aunts age >>>> x=40+y West age>>>> x-14=60-y adding these two eq, gives 2x-14=100 >> 2x=100+14 >> 2x=114 dividing by 2 >>>>>>>>>>> x=57 so the age of aunt is 57 years which is 17 years above 40. and age of west is >> x-14=43 >> 43 years which is 17 years below 60.

OpenStudy (anonymous):

Click "Best Response" to give a medal.